Iodine atoms will combine to form I2 in liquid hexane solvent with a rate constant of 1.5×1010L/mol⋅s. The reaction is second order in I . Since the reaction occurs so quickly, the only way to study the reaction is to create iodine atoms almost instantaneously, usually by photochemical decomposition of I2. Suppose a flash of light creates an initial [I] concentration of 1.00×10−2M .

Question: How long will it take for 94% of the newly created iodine atoms to recombine to form I2? Express your answer using two significant figures.
Answer #1

[I0] = 1*10^-2 M
94% reacted,
so, remaining is 6%
[I] = 6% of 1*10^-2 = 6*10^-4 M


Integrated rate law for 2nd order reation is:
1/[I] = 1/[Io] + k*t
1/(6*10^-4) = 1/(1*10^-2) + (1.5*10^10 *t)
(1.5*10^10 *t) =1566.67
t=1.0*10^-7 s

Answer =1.0*10^-7 s